OpenStudy (anonymous):

. Evaluate: 9C4 126 36 13 20

OpenStudy (lgbasallote):

hint: \[nCr \implies \frac{n!}{r!(n-r)!}\] so \[9C4 \implies \frac{9!}{4!(9-4)!}\] does that help?

OpenStudy (anonymous):

whats the choice

mathslover (mathslover):

what is 9! @black12 ?

mathslover (mathslover):

just tell me in expanded form don't calculate that

OpenStudy (anonymous):

would it be 36

mathslover (mathslover):

hmn i think you are slow with your basics in this topic: \[\large{\color{red}{n! = 1*2*3*4* .... *(n-1)*n}}\] so : \[\large{\color{violet}{9!=1*2*3*4*5*6*7*8*9}}\] got my point?

OpenStudy (anonymous):

so what would the final answer be then

mathslover (mathslover):

sorry we don't provide direct answers, but if you want step by step understanding and soln I can help you with this

mathslover (mathslover):

\[\large{\frac{9!}{4!(9-4)!}=\color{violetred}{\frac{9!}{4!*5!}}=\color{red}{\frac{1*2*3*4*5*6*7*8*9}{1*2*3*4*(1*2*3*4*5)}}}\] can u continue further?

OpenStudy (anonymous):

I got the 9/20 part so it would be 20 then right

mathslover (mathslover):

No no .. you cant multiply 4! * 5! = 20 that is purely wrong you know about factorial?

OpenStudy (anonymous):

if u dont multiply then what?

mathslover (mathslover):

expand 4! 4! = 1*2*3*4 got it?

OpenStudy (anonymous):

ok

mathslover (mathslover):

\[\large{\frac{9!}{4!(9-4)!}=\color{violetred}{\frac{9!}{4!*5!}}=\color{red}{\frac{1*2*3*4*5*6*7*8*9}{1*2*3*4*(1*2*3*4*5)}}}\] now can u tell me what will you get after above multiplications?

OpenStudy (anonymous):

multiply each one of those numbers

mathslover (mathslover):

cancel the terms that are same and then multiply

OpenStudy (anonymous):

this stuff is confusing can we just skip the teaching part. whats the final

mathslover (mathslover):

sorry friend i can not do that.. direct answer not allowed here

OpenStudy (callisto):

Two ways to do it. 1. Use calculator. 2. As the above mentioned, use \[_nC_r = \frac{n!}{r!(n-r)!}\] Which part are you confused at?

OpenStudy (anonymous):

tired too use a calulator and got some big number

OpenStudy (callisto):

Hmm.. Actually, using calculator is faster.. But if you want to do it on your own, it's okay. So, what have you got so far? Which part are you confused at?

OpenStudy (anonymous):

so do i jus calculate all the top numbers toghter then the bottom numbers and reduce

OpenStudy (callisto):

Hmm.. For '!' , you need to 'expand', and then cancel the common factors. \[_9C_4 = \frac{9!}{4!(9-4)!}\]\[_9C_4 = \frac{9\times 8 \times 7 \times 6 \times 5!}{4!(5)!}\] So, can you cancel the common factor here first?

OpenStudy (anonymous):

can u say that again your message looked like this on my screen Hmm.. For '!' , you need to 'expand', and then cancel the common factors. \[_9C_4 = \frac{9!}{4!(9-4)!}\]\[_9C_4 = \frac{9\times 8 \times 7 \times 6 \times 5!}{4!(5)!}\] So, can you cancel the common factor here first

OpenStudy (callisto):

That's it!

OpenStudy (anonymous):

i dont understand

OpenStudy (callisto):

Which part?

OpenStudy (anonymous):

so would it be 126 then

OpenStudy (callisto):

Indeed it is!

OpenStudy (anonymous):